# TAUE: Training-free Noise Transplant and Cultivation Diffusion Model

### v1

@@ 3.2 Composite Generation with Noise Transplantation and Cultivation (NTC)
where $\mathcal{G}_{\sigma}$ is Gaussian blur operator with parameter $\sigma$, and $L_{\text{fg}}^{(1)}$ and $L_{\text{fg}}^{(2)}$ denote the 1st and 2nd channels of the seedling latent $L_{\text{fg}}$ after foreground generation. In parallel, cross-attention maps derived from the foreground prompt $T_{\text{fg}}$ highlight semantically relevant spatial regions. Combining these cues, we define the binary object mask $m_{\text{obj}}$ [^1] The mask is primarily based on attention; minor processing variations depending on usage context are omitted here for clarity. as:
@@
Cross-Attention Blending   To achieve semantic coherence between the foreground and background, we modulate the cross-attention layers using the object mask $m_{\text{obj}}$. As shown in Fig. 3, the foreground prompt $T_{\text{fg}}$ is applied only to the object region, and the background prompt $T_{\text{fg}}$ to the rest. We obtain the mixed cross-attention tensor $A_{\text{mix}}\in\mathbb{R}^{H/8\times W/8\times d}$ by a pixel-wise convex combination of the foreground‑conditioned attention $A_{\text{fg}}$ and the background-conditioned attention $A_{\text{bg}}$:
@@
Denoising Process   We initialize the composite generation with a blended latent that preserves object details while introducing new background content:
@@
Figure 4: Qualitative comparison of layer-wise image generation. For each case, we show the foreground, background, and composite image generated by LayerDiffuse (Zhang and Agrawala 2024), Alfie (Quattrini et al. 2024), and our method. TAUE consistently produces spatially aligned and semantically coherent multi-layer outputs, achieving realistic integration of foreground and background without fine-tuning or inpainting.
3.3 Background Generation with Noise Transplantation and Cultivation
In this final phase, we generate the background image $I_{\text{bg}}$ using the seedling latent $L_{\text{bg}}$ from the composite generation and the background prompt $T_{\text{bg}}$. Similar to the previous phase, we transplant the latent into the initial noise and guide the denoising process with a spatial mask.
We define the background region as the complement of the object mask, $1-m_{\text{obj}}$, ensuring that the foreground remains unaffected during the background generation. The initial latent is defined as:
$z_{\text{bg},T}=(1-m_{\text{obj}})\odot(L_{\text{bg}}+\lambda\cdot n_{t_{\text{crop}}})+m_{\text{obj}}\odot z_{T},$ (11)
where $z_{T}\sim\mathcal{N}(0,1)$ and $\lambda$ controls the influence of the foreground. Denoising starts from $z_{\text{bg},T}$ using a noise blending strategy. During early steps $t_{\text{crop}}\leq t$, we fix the background while allowing the foreground to evolve freely:
$n^{\prime}_{t}=\begin{cases}(1-m_{\text{obj}})\odot n_{t_{\text{crop}}}+m_{\text{obj}}\odot n_{t}&\text{if }t_{\text{crop}}\leq t,\\ n_{t}&\text{otherwise}.\end{cases}$ (12)
This strategy enables independent background cultivation while maintaining spatial and semantic coherence with the fixed foreground. The resulting image $I_{\text{bg}}$ exhibits consistent lighting, structure, and alignment with the composite.
@@ 4.1 Experimental Setup
We adopt SDXL (Podell et al. 2023) as LDM for all experiments and generate images at a resolution of $1024\times 1024$ using the EulerDiscrete scheduler (Karras et al. 2022) with 50 denoising steps. The guidance scale is set to 7.5 for foreground generation and 5.0 for all other cases. To extract the intermediate latent for seedling noise $L_{\text{fg}}$ and $L_{\text{bg}}$, we set $r_{\text{crop}}=0.5$, corresponding to the halfway point of the total denoising process. We compare TAUE with representative layer-wise generation baselines, including the fine-tuned method LayerDiffuse (Zhang and Agrawala 2024) [^2] https://github.com/lllyasviel/LayerDiffuse, and the training-free method Alfie (Quattrini et al. 2024) [^3] https://github.com/aimagelab/Alfie, combined with a background generation pipeline based on outpainting and inpainting (Zhang, Rao, and Agrawala 2023). Other existing methods were not included in our comparison, as their model weights or datasets are not publicly available, precluding reproducible evaluation.
@@ 4.2 Dataset and Metrics
We construct a benchmark of 1770 images filtered from the MS-COCO dataset (Lin et al. 2014), considering limitations of existing methods in handling multi-object generation and extremely small objects. Specifically, we exclude samples where iscrowd = true or the bounding box area is less than $0.03$ relative to image size. This ensures each image contains a single, reasonably sized foreground object suitable for fair comparison. To generate separate prompts for foreground $T_{\text{fg}}$ and background $T_{\text{bg}}$, we employ Phi-3 (Abdin et al. 2024), a language model-based prompt captioning tool. To evaluate overall image quality, we use Fréchet Inception Distance (FID). For semantic consistency, we compute CLIP-Text (CLIP-T) and CLIP-Image (CLIP-I) similarity scores (Hessel et al. 2021), which assess alignment with the textual description and visual fidelity, respectively. To further evaluate the disentanglement of foreground and background from complete scenes, we compute PSNR, SSIM, and LPIPS (Zhang et al. 2018) scores between corresponding regions of the composite image and the individually generated foreground and background layers.
@@ 4.3 Qualitative Result
Fig. 4 presents qualitative results of our method and existing methods. Each method outputs foreground, background, and composite images for the same prompt. Alfie with inpainting, a training-free pipeline, generates the background via outpainting followed by inpainting after foreground generation. However, this often causes foreground features to bleed into the background due to mask misalignment and excessive outpainting. As a result, background images frequently contain residual foreground traces, thereby compromising layer independence. LayerDiffuse, a fine-tuned model, produces better separation and fewer artifacts. However, it still suffers from loss of foreground detail and imperfect semantic harmonization. Composite images occasionally exhibit lighting or shadow inconsistencies between layers.
In contrast, our TAUE generates all layers from a shared latent space via NTC, ensuring semantic and structural coherence without requiring fine-tuning. The resulting foreground, background, and composite outputs remain visually consistent, with no conflicts or missing elements. Notably, harmonization emerges naturally, as all layers are derived from successive noise injections, rather than mask-based composition or attention blending.
@@ 4.4 Quantitative Result
As shown in Tab. 2, TAUE achieves the best image quality among training-free methods across all metrics. Our method outperforms the fine-tuned LayerDiffuse in FID and CLIP-S, indicating superior visual fidelity and stronger alignment with textual prompts. While LayerDiffuse slightly leads in CLIP-I, TAUE remains highly competitive overall.
For layer-wise reconstruction, TAUE achieves the highest foreground accuracy across all metrics (PSNR, SSIM, LPIPS), demonstrating that the transplanted seedling noise effectively preserves object details. Background reconstruction scores are slightly lower than those of Alfie with inpainting and LayerDiffuse, both of which benefit from reusing unmasked background pixels—an advantage that artificially boosts their scores. In contrast, TAUE performs denoising entirely from scratch, relying solely on intermediate seedling noise without reusing any pixel values. This increases the difficulty of the task, making it more challenging, and reduces the likelihood of regenerating the foreground object in exactly the same position and structure. Nevertheless, TAUE maintains a strong background quality and overall consistency, confirming that our NTC reliably guides generation without requiring fine-tuning or additional data.
@@ 5 Applications
We demonstrate the utility of TAUE in three applications: layout and size control, multi-object generation with disentangled attributes, and background replacement. These capabilities are essential in creative workflows that demand spatial precision, semantic separation, and iterative editing. As shown in Fig. 5 and Table 2, TAUE produces high-quality, controllable outputs without requiring fine-tuning or additional training data.
@@
Layout and Size Control   Existing layer-wise image generation models often produce centrally aligned, uniformly sized objects due to limited spatial conditioning (Zhang and Agrawala 2024; Quattrini et al. 2024). Controlling object scale via prompts or layout models often lacks consistency and flexibility. TAUE addresses these limitations by injecting user-defined bounding boxes, which specify the foreground object’s location and size. This guides seedling noise transplantation and denoising to generate semantically coherent content within the desired region.
@@
As shown in Tab. 2, the layout-aware TAUE improves FID, CLIP-I, CLIP-S, and foreground reconstruction (PSNR, SSIM, LPIPS). Explicit size control also enhances alignment with object semantics, yielding more accurate object scales. While background reconstruction is slightly lower due to full denoising instead of pixel reuse, this trade-off improves harmonization and realism. These results show that TAUE enables user-driven spatial generation without training, making it practical for layout-aware applications.
@@
Disentangled Multi-Object Generation   Text-to-image models often suffer from attribute entanglement when generating multiple objects in a single prompt like incorrect color or shape assignments. Prior works (Zhang and Agrawala 2024) mitigate this by sequential generation and compositing, but incur high inference cost and blending artifacts.
@@
TAUE introduces an efficient alternative by transplanting seedling noise to multiple spatial locations in the latent space. Each transplanted latent preserves its object semantics, enabling simultaneous generation of multiple, disentangled objects in a single denoising process. As shown in Fig. 5, our method produces multi-object scenes with semantic separation and spatial consistency, without requiring layout models or additional training. This facilitates controllable scene composition at scale, with minimal overhead.
@@
Background Replacement   Conventional diffusion models regenerate the full image to modify the background, which risks altering the foreground due to entangled representations. TAUE resolves this by decoupling foreground and background generation via latent transplantation. Once the foreground object is generated, its seedling noise can be retained and reused to synthesize a new background independently. This ensures a consistent foreground appearance and layout. Additionally, adjusting the transplantation coordinates allows repositioning of the object across backgrounds, enabling interactive layout-aware editing.
@@
Unlike previous methods that rely on handcrafted prompts or model retraining, TAUE enables seamless background editing in a training-free, zero-shot manner. This enables rapid iteration and fine-grained control in creative workflows, such as UI design and ad generation.
@@ 6 Conclusion
We presented TAUE, a novel training-free framework for layer-wise image generation that enables the synthesis of disentangled foreground and background. By introducing the concept of NTC, TAUE modifies the initial noise within the diffusion process to maintain semantic coherence across layers without requiring any fine-tuning or external datasets. Extensive experiments demonstrate that TAUE achieves performance comparable to fine-tuned models while offering greater flexibility and efficiency. Furthermore, we presented practical applications, such as layout and size control, multi-object generation, and background replacement, highlighting TAUE’s potential for real-world creative workflows. We believe that TAUE opens up new avenues for controllable, modular, and accessible image generation in both research and professional domains.
Limitation   In cases requiring high-fidelity foreground preservation, e.g., when the exact shape, color, or pixel-level structure of the foreground must remain unchanged—TAUE may underperform compared to inpainting-based approaches that modify the background while preserving the foreground. Future work should explore ways to better control this behavior, such as selectively freezing foreground features during compositing or introducing constraints to balance semantic adaptation and structural preservation. Addressing this trade-off between harmonization and fidelity will be crucial to expanding TAUE’s applicability in precision-critical tasks.
Appendix A Ablation Study
We analyze the contribution of each component through controlled ablations.
A.1 High-Pass Filtering for Composite Initialization
[…]
A.2 Choice of Crop Ratio
[…]
Figure 6: Ablation of the Laplacian high-pass filter. Each column shows one prompt. Row 1: composite results without the filter sometimes contain spurious duplicates or softened boundaries. Row 2: applying the filter preserves high-frequency cues in the transplanted latent, producing one clean instance with sharper edges and textures. Best viewed zoomed in.
Figure 7: Effect of crop ratio (25%, 50%, 75%). Columns correspond to the ratio at which the seedling latent is extracted from the denoising trajectory. At 25% (early), prompt-irrelevant residual noise appears; at 75% (late), the foreground fails to harmonize with the regenerated background. The mid-point 50% consistently gives the most coherent composite with accurate lighting and depth.
Appendix B Additional Results
Figure 8 provides additional qualitative examples across diverse categories. For each case we show the Foreground Object, the Background, and the final Composite. TAUE maintains semantic and spatial consistency among all three outputs while preserving fine object details and realistic scene context.
Figure 8: More qualitative results. Each triplet shows foreground, background, and composite. Across vehicles, animals, man-made objects, and natural scenes, TAUE produces crisp foregrounds, harmonized backgrounds, and composites with consistent geometry, lighting, and shadows—without fine-tuning or inpainting.

### v2

@@ 3.3.1 Object Region Mask
where $\mathcal{G}_{\sigma}$ is Gaussian blur operator with parameter $\sigma$, and $L_{\text{fg}}^{(1)}$ and $L_{\text{fg}}^{(2)}$ denote the 1st and 2nd channels of the seedling latent $L_{\text{fg}}$ after foreground generation. In parallel, cross-attention maps derived from the foreground prompt $T_{\text{fg}}$ highlight semantically relevant spatial regions. Combining these cues, we define the binary object mask $m_{\text{obj}}$ as:
@@
3.3.2 Cross-Attention Shearing
To achieve semantic coherence between the foreground and background, we modulate the cross-attention layers using the object mask $m_{\text{obj}}$. As shown in Fig. 3, the foreground prompt $T_{\text{fg}}$ is applied only to the object region, and the background prompt $T_{\text{fg}}$ to the rest. We obtain the mixed cross-attention tensor $A_{\text{mix}}\in\mathbb{R}^{H/8\times W/8\times d}$ by a pixel-wise convex combination of the foreground‑conditioned attention $A_{\text{fg}}$ and the background-conditioned attention $A_{\text{bg}}$:
@@ 3.3.2 Cross-Attention Shearing
3.3.3 Noise Transplant and Cultivation
We initialize the composite generation with a blended latent that preserves object details while introducing new background content:
@@ 3.3.3 Noise Transplant and Cultivation
3.4 Background Generation
This phase mirrors the composite generation but with the object mask inverted. Specifically, we transplant the background latent $L_{\text{bg}}$ into the complementary region $(1-m_{\text{obj}})$ and apply the same noise control scheme symmetrically to cultivate a coherent background $I_{\text{bg}}$ consistent with the composite layers. Unlike the composite phase, we release the mask constraint in the attention layers and apply the background cross-attention $A_{\text{bg}}$ to all spatial positions, allowing the background prompt to globally refine lighting, color, and contextual harmony across the entire scene.
@@ 4.1 Experimental Setup
We adopt SDXL [30] as LDM for all experiments and generate images at a resolution of $1024\times 1024$ using the EulerDiscrete scheduler [23] with 50 denoising steps. The guidance scale is set to 7.5 for foreground generation and 5.0 for all other cases. To extract the intermediate latent for seedling noise $L_{\text{fg}}$ and $L_{\text{bg}}$, we set $r_{\text{crop}}=0.5$, corresponding to the halfway point of the total denoising process. We compare TAUE with representative layer-wise generation baselines, including the fine-tuned method LayerDiffuse [44] [^2] https://github.com/lllyasviel/LayerDiffuse, and the training-free method Alfie [33] [^3] https://github.com/aimagelab/Alfie, combined with a background generation pipeline based on outpainting and inpainting [45]. Other existing methods were not included in our comparison, as their model weights or datasets are not publicly available, precluding reproducible evaluation.
@@ 4.2 Dataset and Metrics
We construct a benchmark of 1,770 images filtered from the MS-COCO dataset [24], considering limitations of existing methods in handling multi-object generation and extremely small objects. Specifically, we exclude samples where iscrowd = true or the bounding box area is less than $0.03$ relative to image size. This ensures each image contains a single, reasonably sized foreground object suitable for fair comparison. To generate separate prompts for foreground $T_{\text{fg}}$ and background $T_{\text{bg}}$, we employ Phi-3 [1], a language model-based prompt captioning tool. To evaluate overall image quality, we use Fréchet Inception Distance (FID). For semantic consistency, we compute CLIP-Text (CLIP-T) and CLIP-Image (CLIP-I) similarity scores [15], which assess alignment with the textual description and visual fidelity, respectively. To further evaluate the disentanglement of foreground and background from complete scenes, we compute PSNR, SSIM, and LPIPS [46] scores between corresponding regions of the composite image and the individually foreground and background layers.
@@ 4.3 Qualitative Result
Fig. 4 presents qualitative results of our method and existing methods. Each method outputs foreground, background, and composite images for the same prompt. Alfie with inpainting, a training-free pipeline, generates the background via outpainting followed by inpainting after foreground generation. However, this often causes foreground features to bleed into the background due to mask misalignment and excessive outpainting. As a result, background images frequently contain residual foreground traces, thereby compromising layer independence. LayerDiffuse, a fine-tuned model, produces better separation and fewer artifacts. However, it still suffers from foreground detail and imperfect semantic harmonization loss. Composite images occasionally exhibit lighting or shadow inconsistencies between layers.
In contrast, our TAUE generates all layers from a shared latent space and attention, ensuring semantic and structural coherence without requiring fine-tuning. The resulting foreground, background, and composite outputs remain visually consistent, with no conflicts or missing elements.
Table 3: Ablation on high-pass filtering and crop ratio. Our default (50% + High-pass) achieves the best holistic quality (lower FID, higher CLIP) while keeping strong reconstruction. Removing the high-pass slightly improves some reconstruction metrics but visibly hurts perceptual quality (edge softness / occasional duplications). Extracting the seedling too late (75%) boosts reconstruction scores but degrades harmonization; too early (25%) leaves residual noise and lowers text alignment.
Overall Quality Layer-wise Reconstruction Quality
Method FID $\downarrow$ CLIP-I $\uparrow$ CLIP-S $\uparrow$ $_{\text{fg}}\uparrow PSNR$$_{\text{bg}}\uparrow$ $_{\text{fg}}\uparrow SSIM$$_{\text{bg}}\uparrow$ $_{\text{fg}}\downarrow LPIPS$$_{\text{bg}}\downarrow$
Ours (50% + High-pass) 55.59 0.655 0.329 23.82 23.55 0.969 0.863 0.045 0.138
50%   w/o High-pass 55.79 0.654 0.328 23.92 23.59 0.970 0.862 0.045 0.139
75% (late) 56.48 0.653 0.328 24.33 25.02 0.974 0.904 0.041 0.091
25% (early) 55.70 0.640 0.321 21.12 19.70 0.953 0.750 0.059 0.284
Figure 5: Ablation on high-pass filtering and crop ratio. Left: removing the Laplacian filter leads to faded or floating objects, while applying it preserves high-frequency details and yields sharper, coherent composites. Right: varying the crop ratio (25%, 50%, 75%) controls when the seedling latent is extracted—25% fail to reconstruct, 75% overfit to the foreground, and 50% achieves the best balance with consistent foreground–background integration.
@@ 4.4 Quantitative Result
As shown in Tab. 2, TAUE achieves the best image quality among training-free methods across all metrics. Our method outperforms the fine-tuned LayerDiffuse in FID and CLIP-S, indicating superior visual fidelity and stronger alignment with the text. While LayerDiffuse slightly leads in CLIP-I, TAUE remains highly competitive overall. For layer-wise reconstruction, TAUE achieves the highest foreground accuracy across all metrics (PSNR, SSIM, LPIPS), demonstrating that the transplanted seedling noise effectively preserves object details. Background reconstruction scores are slightly lower than those of Alfie with inpainting and LayerDiffuse, both of which benefit from reusing unmasked background pixels—an advantage that artificially boosts their scores.
In contrast, TAUE performs denoising entirely from scratch, relying solely on intermediate seedling noise without reusing any pixel values. This increases the difficulty of the task, making it more challenging, and reduces the likelihood of regenerating the foreground object in precisely the same position and structure. Nevertheless, TAUE maintains a strong background quality and overall consistency.
@@
4.5 Ablation Study
As shown in Tab. 3 and Fig. 5, we analyze the effect of the Laplacian high-pass filter and the crop ratio used to extract the seedling latent. Further ablation studies are provided in the supplementary material.
Laplacian high-pass filter Removing the high-pass filter slightly improves reconstruction metrics but degrades perceptual and image quality. The absence of the filter prevents the model from retaining foreground-aware structural information, resulting in artifacts such as faded facial regions in the cat example or floating motorcycles detached from the ground. In contrast, applying the filter preserves high-frequency cues in the transplanted latent, producing sharper edges and visually coherent composites.
Choice of Crop Ratio We also evaluate different crop ratios (25%, 50%, 75%) for latent extraction. Early extraction (25%) fails to capture sufficient foreground structure, often producing incorrect object shapes, whereas late extraction (75%) overemphasizes reconstruction fidelity, yielding floating or scene-inconsistent objects similar to the original foreground. The mid-point setting (50%) provides the best trade-off between structural preservation and generative flexibility, achieving the most coherent composites with balanced foreground–background integration.
@@ 5 Applications
We demonstrate TAUE’s utility in three applications: layout and size control, multi-object generation with disentangled attributes, and background replacement without fine-tuning and special data in Fig. 6 and Tab 2.
@@
Layout and Size Control   Existing layer-wise image generation models often produce centrally aligned, uniformly sized objects due to limited spatial conditioning [44, 33]. Controlling object scale via prompts or layout models often lacks consistency and flexibility. TAUE addresses these limitations by injecting user-defined bounding boxes, which specify the foreground object’s position and size. It guides seedling noise transplantation and denoising to generate semantically coherent content within the desired region.
@@
As shown in Tab. 2, the layout-aware TAUE improves FID, CLIP-I, CLIP-S, and foreground reconstruction (PSNR, SSIM, LPIPS). Explicit size control also enhances alignment with object semantics, yielding more accurate object scales. While background reconstruction is slightly lower due to full denoising instead of pixel reuse, this trade-off improves harmonization and realism.
@@
Disentangled Multi-Object Generation   Text-to-image models often suffer from attribute entanglement when generating multiple objects in a single prompt like incorrect color or shape assignments [5]. Prior works [44] mitigate this by sequential generation and compositing, but incur high inference cost and blending artifacts.
@@
TAUE introduces an efficient alternative by transplanting seedling noise to multiple spatial locations in the latent space. Each transplanted latent preserves its object semantics, enabling simultaneous generation of multiple, disentangled objects in a single denoising process. (Fig. 6,)
@@
Background Replacement   Conventional diffusion models regenerate the full image to modify the background, which risks altering the foreground due to entangled representations. TAUE resolves this by decoupling foreground and background generation via latent transplantation. Once the foreground object is generated, its seedling noise can be retained and reused to synthesize a new background independently. This ensures a consistent foreground appearance and layout. Additionally, adjusting transplantation coordinates allows repositioning of the object across backgrounds, enabling interactive layout-aware editing.
@@
Unlike existing methods that rely on handcrafted prompts or retraining, TAUE enables seamless background edits. This enables rapid iteration and fine-grained control in creative workflows, such as UI design and ad generation.
6 Limitation
In cases requiring high-fidelity foreground preservation, e.g., when the exact shape, color, or pixel-level structure of the foreground must remain unchanged, TAUE may underperform compared to inpainting-based methods that modify the background while preserving the foreground. Future work should explore ways to control this behavior, such as selectively freezing foreground features during compositing or introducing constraints to balance semantic adaptation and structural preservation. Addressing this trade-off between harmonization and fidelity will be crucial to expanding TAUE’s applicability in precision-critical tasks.
@@ 7 Conclusion
We introduced TAUE, a training-free framework that enables layer-wise image generation directly within the latent space of diffusion models, allowing structural and semantic information to propagate across layers with minimal computational overhead. Experimental results show that TAUE matches or surpasses fine-tuned models in both qualitative and quantitative evaluations. Beyond quantitative performance, our framework supports controllable and practical applications such as layout-guided synthesis, multi-object composition, and background replacement, bridging the gap between generative research and real-world creative workflows. We believe TAUE represents a step toward modular, accessible, and layer-aware image generation, opening new directions for diffusion-based content creation and editing.